please help me out with this thank u
Identify the horizontal asymptotes, if any, of the function.
f(x)=1/(x-5)

x = 0
y = 0
x = -5
y = -5
none


Question 2
Identify the vertical asymptotes, if any, of the function.
f(x)=1/(x-5)
x = 0
y = 0
x = -5
y = -5
none


Question 3
The variable z varies jointly with the product of x and y. Use the given values to find an equation that relates the variables, given x=1, y=1/6, z=3.




Question 4
Perform the indicated operation and simplify.
(4x^2-4x/x^2+2x-3)/(4x./x^2+x-6)



Question 5
Perform the indicated operation and simplify.
(4x+1/x^2-4)-(3/x-2)



Question 6
Perform the indicated operation and simplify.
(5/x+1)+(x-7/x-2)


Answer by stanbon(57203)

Identify the horizontal asymptotes, if any, of the function.
f(x)=1/(x-5)
----
As x gets larger and larger f(x) gets closer and closer to 1/x = 0
Horizontal asymptote is y=0
============
Question 2
Identify the vertical asymptotes, if any, of the function.
f(x)=1/(x-5)
Vertical asymptote when denominator = 0.
Verical asymptote at x=5
==============
Question 3
The variable z varies jointly with the product of x and y. Use the given values to find an equation that relates the variables, given x=1, y=1/6, z=3.
z=k(xy)
3 = k(1*1/6)
3 = k(1/6)
k=18
y=18(xy)
==========
Question 4
Perform the indicated operation and simplify.
(4x^2-4x/x^2+2x-3)/(4x./x^2+x-6)
[4x(x-1)/(x+3)(x-1)] / [4x/(x+3)(x-2)]
Cancel where you can to get:
[4x/(x+3)] * [(x+3)(x-2)/4x]
Cancel again to get:
x-2
=========
Question 5
Perform the indicated operation and simplify.
(4x+1/x^2-4)-(3/x-2)
lcm = x^2-4
(4x+1)/lcm -3(x+2)/lcm
[4x+1-3x-2]/lcm
(x-1)/(x^2-4)
==============
Question 6
Perform the indicated operation and simplify.
(5/x+1)+(x-7/x-2)
lcm= (x+1)(x-2)
5(x-2)/lcm + (x-7)(x+1)/lcm
[5x-10+x^2-6x-7]/lcm
[x^2-x-17]/lcm
=========
Cheers,
Stan H.